Circular No. 6007

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994T IN ANONYMOUS GALAXY J. Peters and P. Challis, Harvard-Smithsonian Center for Astrophysics,

report the visual discovery of a supernova (V ∼ 16.0) on June 11.20 UT. Observations were made with the F. L. Whipple Observatory’s 1.5-m Tillinghast telescope (+ FAST spectrograph) and 1.2-m telescopes, yielding a calibrated spectrum that indicates the object to be a type-Ia supernova near maximum brightness. SN 1994T is located at α = 13^{h}16^{m}25^{s}.29, δ = −1^{o}53′16′′.6 (equinox 1950.0), which is 0′′.7 east and 11′′.4 south of the center of the host galaxy.

SUPERNOVA 1994S IN NGC 4495 S. Benetti, European Southern Observatory (ESO); and R. Barbon,

Asiago Astrophysical Observatory, report: “Inspection of a preliminarilyreduced CCD spectrum (range 380–710 nm, resolution 0.8 nm), obtained on June 10.02 UT with the ESO 3.6-m telescope (+ EFOSC1), confirms that SN 1994S is of type Ia, a few days before maximum. The spectrum is dominated by strong P-Cyg lines of intermediate-mass elements: mostly Ca II (H and K), Si II (505.1 and 635.5 nm), Mg II (448.1 nm), and S II (546.8, 561.2, and 565.4 nm). The Si II 597.2-nm line is unusually weak. The expansion velocity deduced from the minimum of the Si II 635.5-nm line, corrected by the recession velocity of 4600 km/s of the parent galaxy (measured from the Na I D interstellar lines), is ∼ 10 400 km/s.”

COMET MUELLER (1993a)

Extension to the ephemeris on IAUC 5964:

1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 17 23^{h}01.^{m}59 −21^{o}43.^{′}0 2.204 2.684 107.^{o}0 21.^{o}2 10.5
27 22 51.61 −26 57.2 2.114 2.763 119.8 18.6 10.5
July 7 22 36.77 −32 38.4 2.058 2.844 132.3 15.3 10.6
17 22 16.45 −38 22.4 2.048 2.926 143.3 12.0 10.7
27 21 50.77 −43 36.8 2.088 3.009 149.7  9.8 10.9
Aug. 6 21 21.08 −47 51.8 2.181 3.093 148.7  9.8 11.1
16 20 50.13 −50 51.2 2.322 3.178 141.6 11.4 11.4
26 20 21.24 −52 38.2 2.503 3.264 131.7 13.4 11.6

1994 June 11 Daniel W. E. Green